OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 4, -4, 4, -4, ...

OpenStudy (anonymous):

@andrewhaze @mathstudent55 @bohotness @undeadknight26

OpenStudy (anonymous):

@campbell_st @Nnesha

OpenStudy (freckles):

it is alternating in sign only hint: think about (-1) as the common ratio.

OpenStudy (bohotness):

what is the question

OpenStudy (anonymous):

find the explicit rule @bohotness

OpenStudy (campbell_st):

well what is each term being multiplied by..?

OpenStudy (campbell_st):

1st term is 4

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

My answer choices are: 1.an=4*1^n-1 2.an=4*1^n+1 3.an=4*(-1)^n 4. an=4*(-1)^n-1

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@skate-4-life270

Nnesha (nnesha):

oky so do you know explicit formula ??

OpenStudy (anonymous):

its positive then negative negative but the difference between the two is 8

OpenStudy (anonymous):

@skate-4-life270 its actually positive negative positive negative....

Nnesha (nnesha):

is it geometric or arithmetic ???

Nnesha (nnesha):

do you know ^^^this ??

OpenStudy (anonymous):

IDK....honestly @Nnesha

Nnesha (nnesha):

okay when number increasing by multiplying or dividing some number then geometric and if you are adding or subtracting previous term to get next one then its arithmetic first terms is 4 2nd one is -4 what you should with 4 to get "negative" 4??

OpenStudy (anonymous):

multiply by -1 @Nnesha

Nnesha (nnesha):

yes right so then its geometric or arithmetic ??

OpenStudy (anonymous):

geometric @Nnesha

Nnesha (nnesha):

yes right

Nnesha (nnesha):

so explicit formula for geometric is \[a _{n} =a _{1} \times r ^{n-1}\]

OpenStudy (anonymous):

The answer is 4 x -1^n so answer 3.

Nnesha (nnesha):

r is ratio a1 is first term

Nnesha (nnesha):

welcome to open study @itsreallykalen http://openstudy.com/code-of-conduct <--- this is for you

OpenStudy (anonymous):

so which one would it be 3 or 4?? @Nnesha @itsreallykalen Because y'all are giving me two different answers

Nnesha (nnesha):

ratios is -1 a1 terms is 4 now replace a1 and r by their values and see which one is ur answer

OpenStudy (anonymous):

so it would be answer 3?? @Nnesha

Nnesha (nnesha):

you mean answer choices= 3

OpenStudy (anonymous):

yes sorry...

Nnesha (nnesha):

nvm

Nnesha (nnesha):

oaky nope its not 3 sorry :(

OpenStudy (anonymous):

So then it would be 4??

Nnesha (nnesha):

 i gave you the formula